Here are summary statistics for randomly selected weights of newborn​ girls: n equals = 186 186​, x overbar x equals = 27.7 27.7 ​hg, s equals = 6.2 6.2 hg. Construct a confidence interval estimate of the mean. Use a 95 95​% confidence level. Are these results very different from the confidence interval 26.2 26.2 hg less than < mu μ less than < 28.2 28.2 hg with only 16 16 sample​ values, x overbar
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 27.7

sample standard deviation = s = 6.2

sample size = n = 186

Degrees of freedom = df = n - 1 = 186 -1 = 185

At 95% confidence level

= 1-0.95% =1-0.95 =0.05

/2 =0.05/ 2= 0.025

t/2,df = t0.025,185 = 1.97

t /2,df = 1.97

Margin of error = E = t/2,df * (s /n)

=1.97 * (6.2 / 186)

Margin of error = E = 0.897

The 95% confidence interval estimate of the population mean is,

- E < <  + E

27.7 - 0.897 < < 27.7 + 0.897

26.303 < < 28.097

(26.303,28.097)